OpenStudy (anonymous):

m2−4m−21m2−5m−14−m2+5m+6m2+2m−3

OpenStudy (anonymous):

@mihirb @agent0smith

OpenStudy (anonymous):

lemme re write it ughhhh

OpenStudy (anonymous):

\[\frac{ m^2 - 4m - 21 }{m^2 -5m - 14 } - \frac{ m^2 + 5m + 6 }{ m^2 + 2m - 3 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me factor it:

OpenStudy (anonymous):

@hartnn oh hey ive missed ya !

hartnn (hartnn):

lol hieee :)

OpenStudy (anonymous):

@mihirb omg whyyyyyy

hartnn (hartnn):

there are 4 quadratics, could you factor any one of them ?

OpenStudy (anonymous):

yes im bout to factor them right now , hold up

OpenStudy (agent0smith):

I'm pretty sure you're able to do this now, just make sure you factor, find the common denominator, and distribute the negative :)

OpenStudy (anonymous):

lol the darn negative

OpenStudy (anonymous):

\[\frac{ ( m - 7 ) ( m + 3 ) }{ ( m + 2 ) ( m - 7 ) } + \frac{ ( m + 3 ) ( m + 2 ) }{ ( m + 3 ) ( m - 1 )}\]

OpenStudy (anonymous):

@agent0smith no its a positive this time

OpenStudy (agent0smith):

Sign in the middle was a minus... oh

OpenStudy (agent0smith):

It's a minus above.

hartnn (hartnn):

means the question had a positive sign ?

OpenStudy (anonymous):

my fault , its a + you guys

hartnn (hartnn):

no problem, go ahead and cancel alll the terms that you can

OpenStudy (anonymous):

\[\frac{ ( m + 3 ) ( m - 1 ) + ( m + 2 ) ( m + 2 ) }{ ( m + 2 ) ( m - 1 )}\]

hartnn (hartnn):

wow, you really need us !? :P

OpenStudy (agent0smith):

Looks right.

OpenStudy (agent0smith):

I said she'd know how to do it! :D

hartnn (hartnn):

looks like agent taught you well :)

OpenStudy (anonymous):

wait im not done lol

hartnn (hartnn):

ofcourse, but there's nothing much to do simplify the numerator

OpenStudy (anonymous):

i have to factor the numerator again

hartnn (hartnn):

what did u get as numerator ?

OpenStudy (anonymous):

its up at the top

hartnn (hartnn):

yes, did u simplify it by distibuting terms ?

OpenStudy (anonymous):

\[\frac{ 2m^2 + 6m + 1 }{ ( m + 2 ) ( m - 1 )}\] as my final answer

hartnn (hartnn):

absolutely correct! :)

OpenStudy (anonymous):

yay (: thank you , i got 100 % on my quiz <3

OpenStudy (agent0smith):

Yay :D

hartnn (hartnn):

omg! \(\large Congratulations! \ddot \smile \)

OpenStudy (anonymous):

thank thank you , i owe it all to my tutors haha

OpenStudy (agent0smith):

You learned fast :) And you had most of the required knowledge down already.

OpenStudy (anonymous):

awe thank you (: